The volume of the pressurised compartment of a jet aircraft is 340.If the pressurised system has to be maintained at temperature of 18 and a cabin altitude of 10000ft when the aircraft is flying 50000 with a complete change of air every minute. Calculate the mass and volume of air per second which must be delivered to the pressurised compartment
1
Expert's answer
2019-09-30T09:53:08-0400

"P_1=P_0(\frac{T_0-0.0065h}{T_0})^{5.35};"

"P_1=101325Pa(\frac{291.15K - 0.0065×3048m}{291.15 K})^{5.35} = 69498.2032 Pa;"


"\rho= \frac{P}{RT};"


"\rho= \frac{69498.2032Pa}{287Nm/kgK×290K};"


"m={\rho}V = 283.9 kg;"

"m = \frac{283.9 kg}{60s }= 4.732 kg/s;"